# Specialist vision-language models for clinical ophthalmology

### v1

@@ Abstract
Clinicians spend a significant amount of time reviewing medical images and transcribing their findings regarding patient diagnosis, referral and treatment in text form. Vision-language models (VLMs), which automatically interpret images and summarize their findings as text, have enormous potential to alleviate clinical workloads and increase patient access to high-quality medical care. While foundational models have stirred considerable interest in the medical community, it is unclear whether their general capabilities translate to real-world clinical utility. In this work, we show that foundation VLMs markedly underperform compared to practicing ophthalmologists on specialist tasks crucial to the care of patients with age-related macular degeneration (AMD). To address this, we initially identified the essential capabilities required for image-based clinical decision-making, and then developed a curriculum to selectively train VLMs in these skills. The resulting model, RetinaVLM, can be instructed to write reports that significantly outperform those written by leading foundation medical VLMs in disease staging (F1 score of 0.63 vs. 0.11) and patient referral (0.67 vs. 0.39), and approaches the diagnostic performance of junior ophthalmologists (who achieve 0.77 and 0.78 on the respective tasks). Furthermore, in a reader study involving two senior ophthalmologists with up to 32 years of experience, RetinaVLM’s reports were found to be similarly correct (78.6% vs. 82.1%) and complete (both 78.6%) as reports written by junior ophthalmologists with up to 10 years of experience. These results demonstrate that our curriculum-based approach provides a blueprint for specializing generalist foundation medical VLMs to handle real-world clinical tasks.
@@ 1 Introduction
Figure 1: We introduce RetinaVLM, the first specialist medical generative vision-language model (VLM). (a) Using a needs-driven training curriculum, we trained RetinaVLM in specialist medical skills that medical foundation VLMs are currently lacking (b) RetinaVLM is able to process retinal optical retinal optical coherence tomography (OCT) images and flexibly respond to text-based queries. (c) Its abilities entail the analysis of imaging biomarkers of age-related macular degeneration (AMD), disease staging, and the referral for treatment.
@@
This current generation of medical language models is fueled by vast amounts of unstructured training data that is extracted from medical textbooks, scientific publications or social media posts of healthcare professionals [9, 6, 10]. These foundation language models have stirred considerable interest among the medical community for their expert-level performance on standardized medical question-answering tasks, such as licensing exams and case studies [12, 13]. However, it is unclear whether this general performance translates to clinical utility in specialist medical domains [14]. Despite its impressive scale, the training data of foundation language models has been collected agnostically towards their downstream application and inherently lacks specialist information related to the challenges faced by clinicians in their daily practice.
In this study, we identify this missing piece in foundation models with the aim of developing generative medical VLMs with real-world clinical utility. We propose to deconstruct clinical problems into sets of mandatory capabilities required for their resolution and selectively train VLMs in these skills. We demonstrate the feasibility of this approach in ophthalmology. To this end we introduce RetinaVLM, the first specialist medical generative VLM (see Figure 1a). RetinaVLM is trained using a dedicated curriculum that is specific to the clinical management of age-related macular degeneration (AMD), the leading cause of blindness in the elderly [15, 16]. It is able to process optical coherence tomography (OCT) images of the retina and flexibly respond to instructions and questions (see Figure 1b). In extensive experiments, we evaluate RetinaVLM’s utility and versatility regarding disease staging, patient referral and biomarker analysis in AMD (see Figure 1c).
@@ 2.1 RetinaVLM, a specialist vision-language model for retinal image analysis
RetinaVLM combines two main components: an ophthalmic vision encoder that processes input OCT images, and a generative LLM that handles textual instructions and outputs the corresponding responses (see Figure 1a). The vision encoder is based on our previous work, in which we found it to perform on par with RETFound [17], a large foundation model for retinal image analysis [18]. We use Meta’s Llama 3 as generative LLM, the best performing model that is openly available at the time of this study [19]. However, without additional training it lacks specialist knowledge related to the analysis of OCT images and the clinical management of AMD. Both these deep neural networks have already been pre-trained on large OCT and natural language datasets, respectively, and we only finetune them in the scope of this study. Additional details about the network architecture and training are included in Section 5.2.
@@
2.2 A curriculum to encode specialist ophthalmological knowledge in vision-language models
An intuitive strategy to specialize VLMs while preserving their ability to flexibly interact with text queries is to provide them with a set of medical images and corresponding question-answer pairs. They are then optimized based on the similarity of their predicted answers to the ground truth. However, visual question-answering (VQA) datasets do not exist for most medical specializations, including ophthalmology.
@@ 2.2 A curriculum to encode specialist ophthalmological knowledge in vision-language models
Together with a large team of ophthalmologists, which are involved with the patient care and academic research of AMD, we defined a set of required capabilities for accurate image-based clinical management of AMD. They include the identification of AMD biomarkers in OCT images, the linking of these to the AMD disease stage, and ultimately deciding on the required referral and treatment of the patient. We then curated a training curriculum, which consists of 41,926 OCT images, and 479,710 visual questions and answers to progressively specialize VLMs in these capabilities.
@@
Figure 2: We curated a two-part curriculum to specialize medical VLMs for clinical use. (a and b) Based on a retrospectively collected OCT imaging dataset, we created a large number of tabular reports as well as a small number of comprehensive textual reports. (c and d) We then used an independent LLM to automatically generate visual question-answers (VQA) based on these reports. (e and f) This yielded two VQA datasets, the first on basic imaging biomarkers of AMD and the second covering more advanced clinical skills. (g and h) Finally, we trained two specialist medical generative VLMs, RetinaVLM-Base and RetinaVLM-Specialist, using either the first or both VQA datasets.
@@
Estimating the disease stage is crucial to patient management as it allows clinicians to monitor and treat patients using standardized protocols. We assessed the ability of four different generative VLMs to determine the AMD disease stage from retinal OCT images. Specifically, we compared two medical foundation VLMs, Med-Flamingo [10] and LLaVA-Med [9], to our two specialist VLMs, RetinaVLM-Base and RetinaVLM-Specialist. Using a testing dataset of previously unseen 276 OCT images, VLMs were tasked to write reports that describe the OCT image before classifying the patient into one of six disease stages (see Figure 3a). The model predictions were compared to ground truth labels obtained from ophthalmologists. Each image was initially graded by two out of six junior ophthalmologists, whose experience in the field ranges from 2 to 15 years. Inter-rater disagreements were resolved by a panel of two senior ophthalmologists with 25 and 32 years of experience, respectively. For additional methodological details, including the instruction given to the VLMs to generate these reports, see Section 5.8.1.
@@
We found that our intermediate RetinaVLM-Base model already performs significantly better than both foundation VLMs, which lack the ophthalmological specialism to stage disease (see Figure 3b). The most performant foundation VLM, Med-Flamingo, achieved a F1 score of 0.11. This was markedly outperformed by our advanced RetinaVLM-Specialist model, scoring at 0.63. This approached, but did not match, the accuracy of the junior ophthalmologists who achieved an F1 score of 0.78. We analyze this discrepancy in detail in Section 3. Both foundation VLMs and RetinaVLM-Base returned a substantial number of invalid reports that did not conclude with one of the six disease stages (see Figure 3c). Conversely, all generated reports by RetinaVLM-Specialist were valid. Similar to human experts, RetinaVLM-Specialist struggled the most when diagnosing wet inactive AMD. We attribute this to the high number of shared imaging biomarkers that indicate either intermediate and late-wet forms of AMD, which sometimes leads to misdiagnosis by both ophthalmologists and RetinaVLM-Specialist (see Figure 3d). This was despite the identification of emerging features related to inactive late wet AMD (a small amount of hyperreflective material, or scar tissue) by RetinaVLM. Four more examples of success and failure cases of RetinaVLM-Specialist are shown in Figure 11a. Moreover, full numerical results as well as the confusion matrix for Med-Flamingo are shown in Figure 12 and 13, respectively.
@@
Figure 4: (a) Correctness, completeness and conciseness of the reports written by LLaVA-Med, RetinaVLM-Specialist and junior ophthalmologists. Reports were scored for on each of the three criteria by senior ophthalmologists using a five-point Likert scale. (b) Representative reports with ratings by one of the senior ophthalmologists.
@@
Eighty-four of the generated reports were scored by the two senior ophthalmologists for their correctness, completeness, and conciseness. They were shown 28 reports written by LLaVA-Med, 28 by RetinaVLM-Specialist, and 28 by the two annotating junior ophthalmologists in random order without knowledge of the author. For each report, the senior ophthalmologist first reviewed the corresponding OCT image before rating the generated report in the three criteria on a five point Likert scale [20].
Echoing our previous findings, the senior ophthalmologists observed that LLaVA-Med failed to comprise factually correct image reports (see Figure 4a), even though it uses specialist terminology that may give the reports the initial appearance of being composed by an ophthalmologist (see Figure 4b). LLaVA-Med also failed to keep reports brief, despite being explicitly instructed to do so.
Conversely, the reports generated by RetinaVLM-Specialist were mostly graded as correct and complete and of similar quality as those written by junior ophthalmologists. Specifically, out of the 28 reports, the senior ophthalmologists either agreed or strongly agreed that 22 (78.6%) written by RetinaVLM-Specialist, and 23 (82.1%) written by junior ophthalmologists, were correct in their observations and conclusions. Similarly, they found that for 20 (71.4%) of the images, neither of these report writers missed any features that were relevant to AMD.
@@
Reports by RetinaVLM-Specialist were deemed to be slightly less concise, where 18 out of 28 (64.3%) were rated as complete compared to 23 (82.1%) by the junior ophthalmologists. This discrepancy can be seen in the second sample in 4b, where RetinaVLM-Specialist correctly identified that the image showed a healthy retina, but also detects a small cyst that was not found in the image, resulting in a lower conciseness rating. Junior ophthalmologists wrote a concise report, but incorrectly associated subretinal drusenoid deposits with intermediate AMD. These shortcomings of both the junior ophthalmologists and RetinaVLM-Specialist are discussed in more detail in Section 3.
@@
We found that both medical foundation VLMs and Retina-Base perform worse than opticians regarding their ability to refer patients in need of urgent treatment (see Figure 5b). While Med-Flamingo failed to refer any of the 29 high-risk patients cases, LLaVA-Med and RetinaVLM-Base were ineffective for differentiating high-risk patients from low- to moderate-risk patients (see Figure 5c). RetinaVLM-Specialist was able to detect 23 out of the 29 high-risk cases that require immediate treatment. At the same time, RetinaVLM’s false discovery rate, defined as the ratio of the number of false positives over the number of predicted positives, of 42.5% is substantially lower than that of opticians at 69.5%. Owing to their ability to better differentiate moderate from high-risk cases, the human ophthalmologists had the lowest false discovery rate of 9.1%, although they simultaneously missed three more cases in urgent need for treatment.

### v2

# Specialized curricula for training vision-language models in retinal image analysis

@@ Abstract
Clinicians spend a significant amount of time reviewing medical images and transcribing their findings regarding patient diagnosis, referral and treatment in text form. Vision-language models (VLMs), which automatically interpret images and summarize their findings as text, have enormous potential to alleviate clinical workloads and increase patient access to high-quality medical care. While foundational models have stirred considerable interest in the medical community, it is unclear whether their general capabilities translate to real-world clinical utility. In this work, we demonstrate that OpenAI’s ChatGPT-4o model, in addition to two foundation VLMs designed for medical use, markedly underperform compared to practicing ophthalmologists on specialist tasks crucial to the care of patients with age-related macular degeneration (AMD). To address this, we initially identified the essential capabilities required for image-based clinical decision-making, and then developed a curriculum to selectively train VLMs in these skills. The resulting model, RetinaVLM, can be instructed to write reports that significantly outperform those written by leading foundation medical VLMs and ChatGPT-4o in disease staging (F1 score of 0.63 vs. 0.33) and patient referral (0.67 vs. 0.50), and approaches the diagnostic performance of junior ophthalmologists (who achieve 0.77 and 0.78 on the respective tasks). Furthermore, in a single-blind reader study two senior ophthalmologists with up to 32 years of experience found RetinaVLM’s reports were found to be substantially more accurate than those by ChatGPT-4o (64.3% vs. 14.3%). These results reinforce that our curriculum-based approach provides a blueprint towards specializing foundation medical VLMs for real-world clinical tasks.
@@ 1 Introduction
Figure 1: We introduce RetinaVLM, a specialist medical generative vision-language model (VLM). (a) Using a curriculum-based approach, we trained RetinaVLM in specialist medical skills that medical foundation VLMs are currently lacking (b) RetinaVLM is able to process retinal optical retinal optical coherence tomography (OCT) images and flexibly respond to text-based queries. (c) Its abilities entail the analysis of imaging biomarkers of age-related macular degeneration (AMD), disease staging, and the referral for treatment.
@@
This current generation of medical language models is fueled by vast amounts of unstructured training data that is extracted from medical textbooks, scientific publications or social media posts of healthcare professionals [9, 6, 10]. These foundation language models have stirred considerable interest among the medical community for their expert-level performance on standardized medical question-answering tasks, such as licensing exams and case studies [12, 13]. However, it is unclear whether this general performance translates to clinical utility in specialist medical domains [14]. Despite its impressive scale, the training data of foundation language models has been collected agnostically of their downstream application. The resulting model is unlikely to acquire the nuanced knowledge necessary for effective application in specialized clinical contexts.
In this study, we identify this missing piece in foundation models towards developing generative medical VLMs with real-world clinical utility. We propose to deconstruct clinical problems into sets of mandatory capabilities required for their resolution and selectively train VLMs in these skills. To train VLMs in these skills, we develop a curriculum-based approach [15] that draws from recent advances in instruction finetuning [16, 17, 18] which iteratively refine models on datasets of increasing quality. We demonstrate the feasibility of this approach in ophthalmology, focussing our analysis to a single retinal disease enables us to assess the clinical limitations of foundation VLMs, and the benefits of training on specialist curricula, in depth. To this end we introduce, RetinaVLM, is a generative medical VLM for OCT images (see Figure 1a). RetinaVLM is trained using a two-part dedicated curriculum that is specific to the clinical management of age-related macular degeneration (AMD), the leading cause of blindness in the elderly [19, 20]. The resulting model is able to process optical coherence tomography (OCT) images of the retina and flexibly respond to instructions and questions (see Figure 1b). In particular, we evaluate RetinaVLM’s utility and versatility regarding disease staging, patient referral and biomarker analysis in AMD (see Figure 1c).
@@ 2.1 RetinaVLM, a specialist vision-language model for retinal image analysis
RetinaVLM combines two main components: an ophthalmic vision encoder that processes input OCT images, and a generative LLM that handles textual instructions and outputs the corresponding responses (see Figure 1a). The vision encoder was trained using self-supervised learning on images from the train set, and was found to perform on par with RETFound [21], a large foundation model for retinal image analysis [22]. For the language model, we use Meta’s Llama 3 as generative LLM which was the most performant, openly available model at the time of this study [23]. Both these deep neural networks have already been pre-trained on large OCT and natural language datasets, respectively. We combine these to create RetinaVLM, following the architectural design of MiniGPT-4 introduced by Zhu et al. [24]. This approach leaves the pretrained models unchanged, and instead trains a third, intermediary network to map visual information from the image encoder to the language model. Additional information regarding the model architecture, training and inference are detailed in Section 5.2.
@@
2.2 A curriculum to encode the capabilities of retinal specialists into vision-language models
An intuitive strategy to specialize VLMs while preserving their ability to flexibly interact with text queries is to provide them with a set of medical images and corresponding visual question-answer (VQA) pairs. VQA-based training, a form of instruction fine-tuning [25, 18], guides VLMs to extract visual information and synthesize it into textual outputs that adhere to specific user directives. This approach also introduces diversity through varying instructions and responses, mitigating overfitting while enhancing the model’s capacity to selectively report relevant clinical features and contextualize findings. However, relevant VQA datasets are scarce for most medical specializations, including ophthalmology.
@@ 2.2 A curriculum to encode the capabilities of retinal specialists into vision-language models
Together with a large team of ophthalmologists, which are involved with the patient care and academic research of AMD, we created a curriculum of VQA datasets designed to train VLMs for assisting image-based clinical decisions regarding AMD. To this end, the ophthalmologists first defined a set of guidelines outlining essential capabilities of agents assisting the image-based clinical management of AMD (verbose versions are documented in Figure 16). Specifically, these include details relating to the identification of AMD biomarkers in OCT images, the linking of these to the AMD disease stage, and ultimately deciding on the required referral and treatment of the patient. These subsequently guided a combination of manual and automated efforts to curate a training curriculum, which consists of 41,926 OCT images, and 479,710 VQA pairs to progressively specialize VLMs in these capabilities.
@@
Figure 2: We curated a two-part curriculum to specialize medical VLMs for clinical use. (a and b) Based on a retrospectively collected OCT imaging dataset, we created a large number of tabular reports as well as a small number of comprehensive textual reports. (c and d) We then used an independent LLM to automatically generate visual question-answers based on these reports. (e and f) This yielded two VQA datasets, the first on basic imaging biomarkers of AMD and the second covering more advanced clinical skills. (g and h) Finally, we trained two specialist medical generative VLMs, RetinaVLM-Base and RetinaVLM-Specialist, using either the first or both VQA datasets.
@@
AMD is a debilitating and irreversible condition marked by the progressive loss of central vision, severely hindering essential activities like reading, driving, and recognizing faces. The demand for timely diagnosis and management currently exceeds the available ophthalmology expertise [26]. Moreover, largely due to aging populations, projections indicate a nearly 50% increase in AMD cases globally to nearly 300 million by 2040 [20]. In this context, automated retinal image analysis emerges as an essential tool to support the interpretation and textual reporting of retinal images.
@@
A key aspect of image report generation involves estimating the disease stage indicated by the retinal image. We assessed the ability of five different generative VLMs to determine the AMD disease stage when writing reports on retinal OCT images. Specifically, we benchmarked two foundation VLMs, Med-Flamingo [10] and LLaVA-Med [9], with purported general abilities in medical image analysis. In addition, we assessed OpenAI’s ChatGPT-4o model [27], developed by OpenAI, L.P. (San Francisco, California, USA). We compared these baseline models against RetinaVLM-Base and RetinaVLM-Specialist, which result from cumulative training on curriculum part 1 and part 2, respectively. Finally we compared these automated models against the overall performance of the six junior ophthalmologists. This experiment was conducted using a testing dataset of 276 previously unseen OCT images, on which VLMs were tasked to write descriptive reports before classifying the patient into one of six disease stages (see Figure 3a). The model predictions were compared to ground truth labels obtained from ophthalmologists. Each image was initially graded by two out of six junior ophthalmologists, who have 2, 3, 5, 8, 10 and 15 years of experience working full-time in ophthalmology clinics after receiving their medical degree. Inter-rater disagreements were resolved by a panel of two senior ophthalmologists with 25 and 32 years of experience, respectively (see Section A.1 for additional details and roles of the junior and senior ophthalmologists). For additional methodological details relating to the the derivation of testing labels and the instruction given to all VLMs to generate these reports, see Sections 5.1.1 and 5.7.1.
@@
We found that both the intermediate RetinaVLM-Base model and ChatGPT-4o perform significantly better than Med-Flamingo and LLaVA-Med, which lack the ophthalmological specialism to stage disease (see Figure 3b). By more effectively classifying conversion to late stages of AMD RetinaVLM-Base and ChatGPT-4o achieve F1 scores of 0.33 and 0.29, respectively. However, beyond this distinction, these models failed to differentiate finer disease stage variations and were markedly outperformed by RetinaVLM-Specialist which scored 0.63 F1. This performance approached, but did not match, the accuracy of the junior ophthalmologists who achieved an F1 score of 0.78. We analyze this last discrepancy in further detail in Section 3. Moreover, foundation VLMs and RetinaVLM-Base returned a substantial number of invalid reports that did not conclude with one of the six disease stages (see Figure 3c). Conversely, all generated reports by RetinaVLM-Specialist were valid. Similar to human experts, RetinaVLM-Specialist struggled the most when diagnosing wet inactive AMD. We attribute this to the high number of shared imaging biomarkers that indicate either intermediate and late-wet forms of AMD, which sometimes leads to misdiagnosis by both ophthalmologists and RetinaVLM-Specialist (see Figure 3d). Four additional examples of success and failure cases of RetinaVLM-Specialist are shown in Figure 11a. Moreover, full numerical results, including a comparison to a standard image-only classification model, as well as the confusion matrix for Med-Flamingo, are shown in Figure 12a and 13, respectively.
@@
Figure 4: (a) Summary statistics of the quality of image reports written by ChatGPT-4o, RetinaVLM-Specialist and junior ophthalmologists, broken down by correctness, completeness and conciseness. Reports were scored for on each of the three criteria by senior ophthalmologists using a five-point Likert scale. (b) Representative reports with ratings by one of the senior ophthalmologists. As ChatGPT-4o tended to write excessively long reports, despite being prompted to shorten them, we display passages the senior ophthalmologists selected as the most important to their given rating. For verbose versions and additional sample reports by ChatGPT-4o see Supplementary Figure 19.
2.3.1 Direct evaluation of imaging reports by senior ophthalmologists
@@
Next, 84 of the generated reports were scored by the two senior ophthalmologists for their correctness, completeness, and conciseness. They were shown 28 reports written by ChatGPT-4o, 28 by RetinaVLM-Specialist, and 28 by the two annotating junior ophthalmologists. We then randomly ordered the 84 reports to decrease the likelihood that multiple reports regarding the same image would appear in succession. Moreover, to mitigate bias we conducted this evaluation as a single-blind study in which the authorship of each report was concealed from the senior ophthalmologists. For each report, the senior ophthalmologist first reviewed the corresponding OCT image before rating the generated report in the three criteria on a five point Likert scale [28].
The senior ophthalmologists observed that ChatGPT-4o largely failed to compose factually correct image reports (see Figure 4a), even though it uses specialist terminology that may give the reports the initial appearance of being written by an ophthalmologist (see Figure 4b). ChatGPT-4o was found to consistently hallucinate the presence of subretinal fluid, and in every relevant instance failed to detect presence of subretinal hyperreflective material and hypertransmission, which are crucial for diagnosing inactive late wet and late dry AMD, respectively. We have included six sample reports by ChatGPT-4o which contain examples of these errors in Supplementary Figure 19, including verbose versions of the three displayed in Figure 4b. Overall, the senior ophthalmologists found that only 4 out of 28 (14.3%) of the reports written by ChatGPT-4o were correct in their observations and conclusions.
Overall senior ophthalmologists either agreed or strongly agreed that reports generated by RetinaVLM-Specialist were more correct (18 vs. 4, or 64.3% vs. 14.3%), complete (18 vs. 5) and concise (18 vs. 4) than those written by ChatGPT-4o. Compared to ChatGPT-4o, RetinaVLM-Specialist correctly detected a wider range of biomarkers, which more frequently led to a correct disease stage estimation. However, there remains a gap in overall performance between RetinaVLM-Specialist and the junior ophthalmologists, with the senior ophthalmologists rating their reports to be more correct (25 vs. 18), complete (24 vs. 18) and concise (22 vs. 18).
@@
An example of this discrepancy can be seen in the second sample in 4b, where RetinaVLM-Specialist correctly identified that the image showed a healthy retina, but also detects a small cyst that was not found in the image. Junior ophthalmologists wrote a concise report, but incorrectly associated subretinal drusenoid deposits with intermediate AMD. Moreover, both RetinaVLM, which has currently only been trained to identify AMD from retinal imaging biomarkers, and ChatGPT-4o fail to report likely cases of retinal vascular disease. These characteristics of both the junior ophthalmologists and RetinaVLM-Specialist are discussed in more detail in Section 3.
@@
We found that both medical foundation VLMs and Retina-Base perform worse than opticians regarding their ability to refer patients in need of urgent treatment (see Figure 5b). While Med-Flamingo failed to refer any of the 29 high-risk patients cases, LLaVA-Med and RetinaVLM-Base were ineffective for differentiating high-risk patients from low- to moderate-risk patients (see Figure 5c). In comparison, ChatGPT-4o was relatively more effective for detecting urgent referrals, but still recommended the referral of 10 patients with low risk, and rarely classified a patient with moderate risk. RetinaVLM-Specialist was the best peforming VLM, and was able to detect 23 out of the 29 high-risk cases that require immediate treatment. At the same time, RetinaVLM’s false discovery rate, defined as the ratio of the number of false positives over the number of predicted positives, of 42.5% is substantially lower than that of opticians at 69.5%. Owing to their ability to better differentiate moderate from high-risk cases, the human ophthalmologists had the lowest false discovery rate of 9.1%, although they simultaneously missed three more cases in urgent need for treatment. A full table of F1 scores for this task, including a comparison to a standard image-only classification model, are shown in Figure 12a.